how many four inch square tiles would be needed to cover a rectangular floor that measures 28 feet by 40 feet
Found 2 solutions by edjones, rfer:
Answer by edjones(8007)

4^2=16 in"
28*12=336
40*12=480
336*480=161280
161280/16=10,080 tiles
.
Ed

Answer by rfer(16322)

28*40=1120 sq ft
1120*9=10,080 tiles